OpenStudy (anonymous):

if a number is increased by 1 then divided by 4, the result is 7

OpenStudy (anonymous):

That number is 27

OpenStudy (anonymous):

(x + 1)/4 = 7 x = (7*4) -1 x = 28

OpenStudy (anonymous):

sorry 27

OpenStudy (anonymous):

Its been 35 years since I have had Algebra . .thanks, I had 28 but still struggling

OpenStudy (anonymous):

(x+1)/4=7 x=(7*4)-1 x=28-1 x=27

OpenStudy (anonymous):

5x2-6=6 - is this equation linear?

OpenStudy (anonymous):

yes the equation is linear